# Protap: A Benchmark for Protein Modeling on Realistic Downstream Applications

### v2

@@ E.1 Model Architecture
$h_{v}^{(i)}$ $\leftarrow\text{LayerNorm}\left(h_{v}^{(i)}+\text{Dropout}(\text{GVP}(h_{v}^{(i)}))\right),$ (3)
@@
ProteinBERT. ProteinBERT [18] is a denoising autoencoder for proteins, inspired by BERT [30] but with a distinct architecture. It takes two inputs: amino acid sequences and GO annotations. The architecture consists of parallel local and global pathways. Local representations are 3D tensors of shape $B\times L\times d_{\text{local}}$ (with $d_{\text{local}}=128$), and global representations are 2D tensors of shape $B\times d_{\text{global}}$ (with $d_{\text{global}}=512$).
Each input sequence is embedded into local features by a shared position-wise embedding layer, while annotations are mapped into global features via a fully connected layer. The model applies six transformer-like blocks [8], each updating local features using both narrow and dilated convolutions (kernel size 9, dilation 1 and 5), followed by feedforward layers. The global path consists of two fully connected layers per block.
Local-global interaction occurs via (1) a broadcast fully connected layer from global to local, and (2) a global attention layer from local to global. Global attention has linear complexity and uses trainable projection matrices $W_{q}$, $W_{k}$, and $W_{v}$ (with $d_{\text{key}}=64$, $d_{\text{value}}=128$), and $n_{\text{heads}}=4$ per block. All activations use GELU [71].
@@
$\phi_{\text{dist}}\!\left(D(i,j)\right)=\Bigl{[}\exp\!\Bigl{(}-\tfrac{(D(i,j)-\mu_{k})^{2}}{2\sigma_{k}^{2}}\Bigr{)}\Bigr{]}_{k=1}^{K},$ (1)
@@
$A_{ij}=\frac{(\mathbf{h}_{i}W_{Q})(\mathbf{h}_{j}W_{K})^{\!\top}}{\sqrt{d}}+b_{ij},$ (2)
@@
$\alpha_{ij}=\operatorname{softmax}_{j}(A_{ij}),\qquad\mathbf{z}_{i}=\sum_{j=1}^{L}\alpha_{ij}(\mathbf{h}_{j}W_{V}),$ (3)
@@
ESM2. ESM-2 [17] is a family of transformer-based protein language models trained to predict masked amino acids in protein sequences. Compared to its predecessor ESM-1b [16], ESM-2 introduces architectural refinements, improved training procedures, and is scaled across model sizes ranging from 8M to 15B parameters. The model is trained using a standard masked language modeling (MLM) objective: for 15% randomly masked positions in a sequence, the model predicts the identity of each masked amino acid based on its unmasked context.
@@
$\mathcal{L}_{\text{MLM}}=\sum_{i\in M}\log p(x_{i}\mid x_{\setminus M}),$ (1)
where $M$ is the set of masked positions and $x_{\setminus M}$ denotes the observed amino acids. Training is performed on $\sim$65 million sequences sampled from $\sim$43 million UniRef50 [42] clusters, covering $\sim$138 million UniRef90 entries.
@@
E.2 Pretrain Model Experimental Setup
@@
Table 17: Hyperparameter setting of the EGNN pre-training.
@@
Table 18: Hyperparameter setting of the SE(3) Transformer pre-training.
@@
Table 19: Hyperparameter setting of the D-Transformer pre-training.
@@
Table 20: Hyperparameter setting of the GVP pre-training.
@@
Table 21: Hyperparameter setting of the ProteinBERT pre-training.
@@
Pre-training Data. The structural information used for pretraining is derived from the AFDB Swiss-Prot dataset, while the functional and family annotations are obtained from UniProt Swiss-Prot. Swiss-Prot is a high-quality, manually curated protein sequence database within UniProt. Its goal is to provide comprehensive and biologically meaningful annotations of protein sequences, such as their functions, families, and maintain minimal redundancy. The AFDB Swiss-Prot dataset is a subset of the AlphaFold Protein Structure Database that specifically contains structure predictions for all protein sequences corresponding to UniProt Swiss-Prot entries. Currently, the dataset includes a total of 542,378 entries.
Settings. We adopt task and architecture-specific training and model hyperparameters across different pretraining tasks and model types. Detailed configurations are summarized in Tab. 17-21.

### v3

@@ G.1. Model Architecture
(3) $h_{v}^{(i)}$ $\leftarrow\text{LayerNorm}\left(h_{v}^{(i)}+\text{Dropout}(\text{GVP}(h_{v}^{(i)}))\right),$
@@
ProteinBERT. ProteinBERT (Brandes et al., 2022) is a denoising autoencoder for proteins, inspired by BERT (Devlin et al., 2019) but with a distinct architecture. It takes two inputs: amino acid sequences and GO annotations. The architecture consists of parallel local and global pathways. Local representations are 3D tensors of shape $B\times L\times d_{\text{local}}$ (with $d_{\text{local}}=128$), and global representations are 2D tensors of shape $B\times d_{\text{global}}$ (with $d_{\text{global}}=512$).
Each input sequence is embedded into local features by a shared position-wise embedding layer, while annotations are mapped into global features via a fully connected layer. The model applies six transformer-like blocks (Vaswani et al., 2017), each updating local features using both narrow and dilated convolutions (kernel size 9, dilation 1 and 5), followed by feedforward layers. The global path consists of two fully connected layers per block.
Local-global interaction occurs via (1) a broadcast fully connected layer from global to local, and (2) a global attention layer from local to global. Global attention has linear complexity and uses trainable projection matrices $W_{q}$, $W_{k}$, and $W_{v}$ (with $d_{\text{key}}=64$, $d_{\text{value}}=128$), and $n_{\text{heads}}=4$ per block. All activations use GELU (Hendrycks and Gimpel, 2016).
@@
(1) $\phi_{\text{dist}}\!\left(D(i,j)\right)=\Bigl[\exp\!\Bigl(-\tfrac{(D(i,j)-\mu_{k})^{2}}{2\sigma_{k}^{2}}\Bigr)\Bigr]_{k=1}^{K},$
@@
(2) $A_{ij}=\frac{(\mathbf{h}_{i}W_{Q})(\mathbf{h}_{j}W_{K})^{\!\top}}{\sqrt{d}}+b_{ij},$
@@
(3) $\alpha_{ij}=\operatorname{softmax}_{j}(A_{ij}),\qquad\mathbf{z}_{i}=\sum_{j=1}^{L}\alpha_{ij}(\mathbf{h}_{j}W_{V}),$
@@
ESM2. ESM-2 (Lin et al., 2023) is a family of transformer-based protein language models trained to predict masked amino acids in protein sequences. Compared to its predecessor ESM-1b (Rives et al., 2021), ESM-2 introduces architectural refinements, improved training procedures, and is scaled across model sizes ranging from 8M to 15B parameters. The model is trained using a standard masked language modeling (MLM) objective: for 15% randomly masked positions in a sequence, the model predicts the identity of each masked amino acid based on its unmasked context.
@@
(1) $\mathcal{L}_{\text{MLM}}=\sum_{i\in M}\log p(x_{i}\mid x_{\setminus M}),$
where $M$ is the set of masked positions and $x_{\setminus M}$ denotes the observed amino acids. Training is performed on $\sim$65 million sequences sampled from $\sim$43 million UniRef50 (Consortium, 2019) clusters, covering $\sim$138 million UniRef90 entries.
@@
G.2. Pretrain Model Experimental Setup
@@
Table 16. Hyperparameter setting of the EGNN pre-training.
@@
Table 17. Hyperparameter setting of the SE(3) Transformer pre-training.
@@
Table 18. Hyperparameter setting of the D-Transformer pre-training.
@@
Table 19. Hyperparameter setting of the GVP pre-training.
@@
Table 20. Hyperparameter setting of the ProteinBERT pre-training.
@@
Pre-training Data. The structural information used for pretraining is derived from the AFDB Swiss-Prot dataset, while the functional and family annotations are obtained from UniProt Swiss-Prot. Swiss-Prot is a high-quality, manually curated protein sequence database within UniProt. Its goal is to provide comprehensive and biologically meaningful annotations of protein sequences, such as their functions, families, and maintain minimal redundancy. The AFDB Swiss-Prot dataset is a subset of the AlphaFold Protein Structure Database that specifically contains structure predictions for all protein sequences corresponding to UniProt Swiss-Prot entries. Currently, the dataset includes a total of 542,378 entries. ESM-2 and SaProt are pretrained on UniRef sequences clustered at the 50% sequence identity level. ESM-C is also pretrained on UniRef but with clustering at the 70% sequence identity level, and further incorporates protein sequences from the MGnify and Joint Genome Institute (JGI) databases, resulting in 83M, 372M, and 2B clusters for UniRef, MGnify, and JGI, respectively.
Settings. We adopt task and architecture-specific training and model hyperparameters across different pretraining tasks and model types. Detailed configurations are summarized in Tab. 16-20.
@@
Appendix H Domain-Specific Model Architecture
[…]
(1) $h^{(\ell)}_{i}=W_{1}^{(\ell)}h^{(\ell-1)}_{i}+\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\left(W_{2}^{(\ell)}h^{(\ell-1)}_{j}+W_{3}^{(\ell)}e_{ij}\right).$
With attention weights defined as:
(2) $\alpha_{ij}=\text{softmax}\left(\frac{\left(W_{4}^{(\ell)}h_{i}^{(\ell-1)}\right)^{\top}\left(W_{5}^{(\ell)}h_{j}^{(\ell-1)}+W_{3}^{(\ell)}e_{ij}\right)}{\sqrt{d_{\ell}}}\right).$
DeepPROTACs. DeepPROTAC (Li et al., 2022) is a deep learning framework designed to predict the degradation efficacy of PROTAC molecules by integrating structural and chemical information from multiple components: the protein of interest (POI), the E3 ligase, and the PROTAC compound itself. The input includes five parts: the POI binding pocket, the E3 ligase binding pocket, the warhead (the PROTAC moiety binding to the POI), the E3 ligand (binding to the E3 ligase), and the linker represented as a SMILES string. For the four molecular structures (POI pocket, E3 pocket, warhead, and E3 ligand), atom-level graphs are constructed and processed using Graph Convolutional Layers (GCLs) followed by max pooling to obtain fixed-size feature embeddings. The linker is embedded using an LSTM network to capture sequential chemical features. All five embeddings are concatenated and passed through fully connected layers to yield a binary output: 1 indicates good degradation (defined as $\mathrm{DC}_{50}\leq 100\,\mathrm{nM}$ and $\mathrm{D}_{\max}\geq 80\%$), while 0 indicates poor degradation ($\mathrm{DC}_{50}>100\,\mathrm{nM}$ or $\mathrm{D}_{\max}<80\%$). This modular architecture enables DeepPROTACs to effectively model ternary complex formation and predict degradation outcomes based on both structural and sequential representations.
ET-PROTACs. ET-PROTACs (Cai et al., 2025) is an end-to-end deep learning model designed for PROTAC degradation prediction, consisting of four main components: (i) initial PROTAC and protein featurization; (ii) representation learning using 3D graph-based and sequence-based encoders; (iii) a cross-modal ternary attention block; and (iv) a final classifier. Each component is described in detail below. Each PROTAC is represented as a 2D molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and associated with a 3D coordinate matrix $X\in\mathbb{R}^{|\mathcal{V}|\times 3}$ using RDKit [^1] https://www.rdkit.org/. Nodes $v_{i}\in\mathcal{V}$ represent atoms, and edges $a_{ij}\in\mathcal{E}$ denote chemical bonds. Node features $h_{i}$ and edge features $a_{ij}$ are encoded using learned embeddings. Additionally, atomic 3D coordinates $x_{i}\in\mathbb{R}^{3}$ are embedded to capture spatial information. The input protein is given as a sequence $P=\{a_{1},a_{2},\ldots,a_{n}\}$ of 23 amino acids (including one non-standard residue). Protein sequences are encoded by two embedding layers: a learned character embedding of dimension 64, and a positional embedding of the same dimension. The molecular graph and 3D coordinates are processed by an Equivariant Graph Neural Network (EGNN), which is invariant to rotations and translations. Let $h_{i}^{\ell}$, $x_{i}^{\ell}$, and $a_{ij}$ be the node features, coordinates, and edge features at layer $\ell$, respectively. The EGNN updates are as follows:
(1) $m_{ij}$ $=\phi_{e}(h_{i}^{\ell},h_{j}^{\ell},\|x_{i}^{\ell}-x_{j}^{\ell}\|^{2},a_{ij}),$
$m_{i}$ $=\sum_{j\in\mathcal{N}(i)}m_{ij},$
$h_{i}^{\ell+1}$ $=\phi_{h}(h_{i}^{\ell},m_{i}),$
where C=1/(M1)C=1/(M-1), $\phi_{e}$, $\phi_{x}$, and $\phi_{h}$ are learnable functions, and $M$ is the number of atoms. Final PROTAC embeddings are obtained by combining features of the warhead, linker, and E3 ligase substructures. ET-PROTACs leverages CNNs to encode protein sequences directly from FASTA format. The sequence embedding $E$ is passed through three CNN layers to yield latent protein features $P_{\text{cnn}}\in\mathbb{R}^{d\times f}$, where $f$ is the number of filters in the last layer. Given feature matrices $D=\{d_{1},\ldots,d_{X}\}$ (PROTAC), $P=\{p_{1},\ldots,p_{Y}\}$ (protein), and $L=\{l_{1},\ldots,l_{Z}\}$ (ligase), we form composite pairs:
$\text{pd}_{i}$ $=\text{CONCAT}(p_{j},d_{i}),$
$\text{dl}_{j}$ $=\text{CONCAT}(d_{i},l_{k}),$
$\text{pda}_{i}$ $=F(W_{\text{pd}}\cdot\text{pd}_{i}+b),$
$\text{dla}_{j}$ $=F(W_{\text{dl}}\cdot\text{dl}_{j}+b),$
$A_{i,j}$ $=F(W_{a}\cdot(\text{pda}_{i}+\text{dla}_{j})+b),$
$A_{\text{pd}}$ $=\sigma\left(\text{MEAN}(A,2)\right),$
$A_{\text{dl}}$ $=\sigma\left(\text{MEAN}(A,1)\right),$
DeepFRI. DeepFRI (Gligorijević et al., 2021) is a deep learning framework for protein function prediction based solely on amino acid sequences. Given a protein sequence of length $L$, the input is encoded as a binary matrix $X=[x_{1},\ldots,x_{L}]\in\{0,1\}^{L\times 26}$, where each $x_{i}$ is a one-hot vector indicating the amino acid type at position $i$, including 20 standard residues, 5 non-standard residues, and a gap symbol. This input is passed through a set of one-dimensional convolutional layers, each consisting of $f_{n}=512$ filters of varying kernel lengths $f_{l}$, followed by rectified linear unit (ReLU) activation, $\text{ReLU}(x)=\max(x,0)$, and a global max-pooling operation. The use of multiple filter sizes enables the extraction of complementary local patterns from the sequence. Outputs from 16 such CNN layers are concatenated, resulting in an $L\times 8192$ feature representation. Finally, a fully connected layer with sigmoid activation outputs probabilities for either Gene Ontology (GO) terms or Enzyme Commission (EC) classes. The output dimensionality is task-specific, corresponding to $|\text{GO}|$ or $|\text{EC}|$ respectively.
DPFunc. DPFunc (Wang et al., 2025) is a structure-aware protein function prediction framework that captures residue-level information by integrating 3D geometric structure and pretrained sequence embeddings. For a given protein of length $l$, a residue-level undirected graph is constructed where each node represents a residue, and an edge is added between two residues if the distance between their Cα atoms is less than 10 Å. This defines the adjacency matrix $A\in\{0,1\}^{l\times l}$. Node features are initialized using embeddings from a pretrained protein language model (ESM-1b), resulting in $X\in\mathbb{R}^{l\times d}$. Two graph convolutional layers are then used to propagate structural information, with residual connections added to facilitate gradient flow. The update rule for the $k$-th GCN layer is given by:
(1) $X^{(k+1)}=X^{(k)}+\text{ReLU}\left(\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}X^{(k)}W^{(k)}\right),$
(2) $H=\text{ReLU}\left(\left(\text{ReLU}(IPR\cdot W_{\text{emd}})W_{1}+b_{1}\right)W_{2}+b_{2}\right).$
(3) $Q_{i}=HW^{Q}_{i},\quad K_{i}=X_{\text{final}}W^{K}_{i},\quad V_{i}=X_{\text{final}}W^{V}_{i},$
and attention weights are derived as:
(4) $W^{A}_{i}=\text{Softmax}\left(\frac{K_{i}Q_{i}^{\top}}{\sqrt{d}}\right).$
The multi-head attention output is computed as:
(4) $X_{\text{multi}}=\text{LayerNorm}\left(\text{Concat}(W^{A}_{1}V_{1},\ldots,W^{A}_{n}V_{n})W_{\text{trans}}+X_{\text{final}}\right),$
followed by a feedforward transformation with residual connection:
(5) $X_{\text{out}}=\text{LayerNorm}\left(\text{FF}(X_{\text{multi}})+X_{\text{multi}}\right).$
(6) $x_{\text{pool}}=\sum_{i=1}^{l}X_{\text{out}}[i].$
(7) $x_{\text{integrate}}=\text{Concat}\left(x_{\text{pool}},\frac{1}{l}\sum_{i=1}^{l}X[i]\right),$
(8) $\hat{y}=\text{Sigmoid}(\text{MLP}(x_{\text{integrate}})).$
(9) $\hat{y}^{\text{post}}_{i}=\max\left(\hat{y}_{i},\hat{y}_{\text{child}_{1}},\ldots,\hat{y}_{\text{child}_{n}}\right).$
UniZyme. UniZyme (Li et al., 2025) is a biochemically informed framework designed to generalize protein cleavage site prediction across diverse enzymes. It comprises an enzyme encoder and a substrate encoder. The enzyme encoder integrates sequence-derived features with energetic frustration and 3D structural information. Given an enzyme $P_{e}=(X,R)$, where $X$ are residue embeddings and $R$ are Cα coordinates, the pairwise frustration score is defined as:
(1) $F(i,j)=\frac{E(i,j)-\mu_{\text{rand}}(i,j)}{\sigma_{\text{rand}}(i,j)}.$
(2) $\Phi^{\text{energy}}_{i,j}=\text{MLP}(\phi_{\text{energy}}(F(i,j))),\quad\Phi^{\text{dist}}_{i,j}=\text{MLP}(\phi_{\text{dist}}(\|r_{i}-r_{j}\|_{2})).$
(3) $A^{k}_{i,j}=\frac{(h^{k-1}_{i}W_{Q})(h^{k-1}_{j}W_{K})^{T}}{\sqrt{d}}+\Phi^{\text{energy}}_{i,j}+\Phi^{\text{dist}}_{i,j}.$
(4) $\hat{a}_{i}=\sigma(h_{i}\cdot w_{a}).$
(5) $h_{e}=\sum_{i=1}^{N}\frac{f(\hat{a}_{i})}{\sum_{j}f(\hat{a}_{j})}h_{i}.$
(6) $\hat{c}_{e,s}^{(t)}=\text{MLP}(\text{CONCAT}(H_{s}^{t:t+l},h_{e})).$
(7) $\mathcal{L}=\mathcal{L}_{c}(D_{c})+\lambda\mathcal{L}_{a}(D_{a}).$
MONN. MONN (Li et al., 2020b) is composed of four interconnected modules: a graph convolutional module for molecular representation, a CNN module for residue-level protein representation, a pairwise interaction module for atom-residue interaction estimation, and an affinity prediction module for compound-protein binding affinity estimation.
(1) $v_{i}^{0}=f(W_{\text{init}}v^{\text{init}}_{i}),$
(2) $t_{i}^{l}=\sum_{v_{k}\in\mathcal{N}(v_{i})}f(W^{l}_{\text{gather}}[v_{k}^{l-1},e_{i,k}]),$
followed by feature updates:
(3) $u_{i}^{l}=f(W^{l}_{\text{update}}[t_{i}^{l},v_{i}^{l-1}]).$
Protein sequences are encoded by mapping residues to BLOSUM62 (Eddy, 2004) columns and processed through a 1D CNN to obtain residue embeddings $\{r_{j}\}$ in $\mathbb{R}^{h_{1}}$.
(4) $P_{i,j}=\sigma(f(W_{\text{atom}}v_{i})\cdot f(W_{\text{residue}}r_{j})).$
(5) $h_{v,i}$ $=f(W_{v}v_{i}),\quad h_{s}=f(W_{s}s),\quad h_{r,j}=f(W_{r}r_{j}),$
(6) $h_{c}$ $=\sum_{i}\alpha_{v,i}h_{v,i},\quad h_{p}=\sum_{j}\alpha_{r,j}h_{r,j}.$
The binding affinity is predicted by a regression on the outer product between $h_{c},h_{s}]$ and $h_{p}$:
(7) $a=W_{\text{affinity}}f(\text{flatten}([h_{c},h_{s}]\otimes h_{p})),$
where $W_{\text{affinity}}\in\mathbb{R}^{1\times 2h_{2}^{2}}$.
CLIPZyme. CLIPZyme (Mikhael et al., 2024) formulates enzyme screening as a retrieval task, where a predefined set of enzymes is ranked based on their predicted ability to catalyze a given chemical reaction. Each reaction $R$ and enzyme $P$ is encoded into a $d$-dimensional vector $r,p\in\mathbb{R}^{d}$ using learned encoders $f_{\text{rxn}}$ and $f_{\text{p}}$, respectively. A scoring function $s(r,p)$ computes the cosine similarity between the two embeddings:
(1) $s_{ij}=s(r_{i},p_{j})=\frac{r_{i}\cdot p_{j}}{\|r_{i}\|\|p_{j}\|}.$
(2) $\mathcal{L}_{ij}=-\frac{1}{2N}\left(\log\frac{e^{s_{ij}/\tau}}{\sum_{k}e^{s_{ik}/\tau}}+\log\frac{e^{s_{ij}/\tau}}{\sum_{k}e^{s_{kj}/\tau}}\right),$
To represent a reaction, atom-mapped molecular graphs of the reactants $G_{x}$ and products $G_{y}$ are first encoded by a directed message passing neural network (Yang et al., 2019) (DMPNN) $f_{\text{mol}}$, yielding atom features $a_{i}$ and bond features $b_{ij}$:
(3) $a_{i},b_{ij}=f_{\text{mol}}(G_{x},G_{y}).$
(3) $v_{i}^{(TS)}$ $=v_{i}^{(x)}=v_{i}^{(y)},$
(4) $e_{ij}^{(TS)}$ $=b_{ij}^{(x)}+b_{ij}^{(y)}.$
(4) $a^{\prime}_{i},b^{\prime}_{ij}$ $=f_{TS}(G_{TS}),$
(5) $r$ $=\sum_{i}a^{\prime}_{i}.$